# Classifying localizing subcategories of a Grothendieck category

### v1

@@ 6. In the case of commutative (coherent) rings
$\operatorname{Spec}A\supseteq\mathcal{V}\stackrel{{\scriptstyle\mathcal{O}}}{{\mapsto}}\mathcal{O}_{\mathcal{V}}=\{\overline{A/\mathfrak{p}}\in\operatorname{ASpec}A|\hskip 2.84544pt\mathfrak{p}\in\mathcal{V}\}\subseteq\operatorname{ASpec}A\hskip 4.26773ptand$
$\operatorname{ASpec}A\supseteq\hskip 2.84544pt\mathcal{O}\stackrel{{\scriptstyle\mathcal{V}}}{{\mapsto}}\mathcal{V}_{\mathcal{O}}=\{\mathfrak{p}\in\operatorname{Spec}A|\hskip 2.84544pt\overline{A/\mathfrak{p}}\in\mathcal{O}\}\subseteq\operatorname{Spec}A$
establish a bijective correspondence between the lattices of specialization-closed subsets of $\operatorname{Spec}A$ and the lattice of open subsets of $\operatorname{ASpec}A$.
Proof.
For every specialization-closed subset $\mathcal{V}$ of $\operatorname{Spec}A$ and $\alpha\in\mathcal{O}_{\mathcal{V}}$, by section 5, there exists $\mathfrak{p}\in\operatorname{Spec}A$ such that $\alpha=\overline{A/\mathfrak{p}}$ and so $\mathfrak{p}\in\mathcal{V}$. Now, for every $\beta\in\operatorname{ASupp}A/\mathfrak{p}$, there exists a prime ideal $\mathfrak{q}$ of $A$ such that $A/\mathfrak{q}$ is a subquotient of $A/\mathfrak{p}$. This implies that $\mathfrak{p}\subset\mathfrak{q}$. Since $\mathcal{V}$ is a specialization-closed subset of $\operatorname{Spec}A$, we deduce that $\mathfrak{q}\in\mathcal{V}$; and hence $\beta\in\mathcal{O}_{\mathcal{V}}$. Therefore $\operatorname{ASupp}A/\mathfrak{p}\subset\mathcal{O}_{\mathcal{V}}$ so that $\mathcal{O}_{\mathcal{V}}$ is an open subset of $\operatorname{ASpec}A$. For any open subset $\mathcal{O}$ of $\operatorname{ASpec}A$ and any prime ideals $\mathfrak{p},\mathfrak{q}$ of $A$ with $\mathfrak{p}\subseteq\mathfrak{q}$ and $\mathfrak{p}\in\mathcal{V}_{\mathcal{O}}$, we have $\overline{A/\mathfrak{p}}\in\mathcal{O}$. This implies that $\operatorname{ASupp}A/\mathfrak{p}\subseteq\mathcal{O}$ as $\mathcal{O}$ is an open subset of $\operatorname{ASpec}A$. The canonical epimorphism $A/\mathfrak{p}\twoheadrightarrow A/\mathfrak{q}$ implies that $\overline{A/\mathfrak{q}}\in\operatorname{ASupp}A/\mathfrak{p}\subset\mathcal{O}$ so that $\mathfrak{q}\in\mathcal{V}_{\mathcal{O}}$. Now, it is easy to see that $\mathcal{O}_{\mathcal{V}_{\mathcal{O}}}=\mathcal{O}$ and $\mathcal{V}_{\mathcal{O}_{\mathcal{V}}}=\mathcal{V}$. ∎
[…]
The set $\{\operatorname{ZSupp}A/\mathfrak{a}|\hskip 2.84544pt\mathfrak{a}$ is an ideal of $\mathcal{A}\}$ forms a basis for the topology $\operatorname{Sp}A$.
Proof.
Any indecomposable injective module $E$ contains a cyclic submodule $A/\mathfrak{c}$. This implies that $E\in\operatorname{ZSupp}A/\mathfrak{c}$. For any ideals $\mathfrak{a}$ and $\mathfrak{b}$ of $A$, it is clear that $\operatorname{ZSupp}A/(\mathfrak{a}+\mathfrak{b})\subseteq\operatorname{ZSupp}A/\mathfrak{a}\cap\operatorname{ZSupp}A/\mathfrak{b}$. The rest of the proof is similar to the proof of section 6. ∎
A set $\mathfrak{F}$ of ideals of $A$ is said to be a Gabriel topology on $A$ is it satisfies:
(2). If $\mathfrak{a}$, $\mathfrak{b}$ belonge to $\mathfrak{F}$, then $\mathfrak{a}\cap\mathfrak{b}\in\mathfrak{F}$.
Let $E$ be an indecomposable injective module in $\operatorname{Sp}A$. Clearly, the Gabriel topology
$\mathfrak{F}_{E}=\{\mathfrak{a}|\hskip 2.84544ptE\notin\operatorname{ZSupp}A/\mathfrak{a}\}$
is associated to the localizing subcategory cogenerated by $E$ as follows:
$\mathcal{X}(E)=\{M\in\operatorname{Mod}A|\hskip 2.84544pt\operatorname{Hom}_{A}(M,E)=0.$
More generally, let $\mathcal{E}$ be a subset of $\operatorname{Sp}A$. To $\mathcal{E}$, we associated a Gabriel topology
$\mathfrak{F}_{\mathcal{E}}=\bigcap_{E\in\mathcal{E}}\mathfrak{F}_{E}=\{\mathfrak{a}|\hskip 2.84544pt(\operatorname{ZSupp}A/\mathfrak{a})\cap\mathcal{E}=\emptyset\}.$
[…]
Given a Gabriel topology $\mathfrak{F}$, an $A$-module $M$ is said to be $\mathfrak{F}$-discrete if $\operatorname{Ann}(x)\in\mathfrak{F}$ for all $x\in M$. The following theorem provides a new classification of Gabriel topologies on $A$. There are many classifications of Gabriel topologies on $A$ (e.g., see [G, St]). In the following theorem, we provide an alternative classification:
[…]
The maps $\mathfrak{F}\stackrel{{\scriptstyle D}}{{\mapsto}}D(\mathfrak{F})$ and $D\stackrel{{\scriptstyle\mathfrak{F}}}{{\mapsto}}\mathfrak{F}_{D}.$ establish a bijective correspondence between the lattices of Gabriel topologies on $A$ and the lattice of closed subsets of $\operatorname{Sp}A$.
Proof.
Let $\mathfrak{F}$ be a Gabriel topology on $A$. We show that $\mathfrak{F}_{D(\mathfrak{F})}=\mathfrak{F}$. Given $\mathfrak{a}\in\mathfrak{F}$, the definition implies that $\operatorname{Hom}_{A}(A/\mathfrak{a},D(\mathfrak{F}))=0$ so that $\mathfrak{a}\in\mathfrak{F}_{D(\mathfrak{F})}$. Conversely, if $\mathfrak{b}\in\mathfrak{F}_{D(\mathfrak{F})}$, then $\operatorname{Hom}_{A}(A/\mathfrak{b},D(\mathfrak{F}))=0$ so that $\operatorname{ZSupp}A/\mathfrak{b}\subseteq\bigcup_{\mathfrak{a}\in\mathfrak{F}}\operatorname{ZSupp}A/\mathfrak{a}$. It then follows from section 3 and section 6 that $A/\mathfrak{b}\in\langle A/\mathfrak{a}|\hskip 2.84544pt\mathfrak{a}\in\mathfrak{F}\rangle_{\rm loc}$, where $\langle A/\mathfrak{a}|\hskip 2.84544pt\mathfrak{a}\in\mathfrak{F}\rangle_{\rm loc}$ is the localizing subcategory of $\operatorname{Mod}A$ associated to $\mathfrak{F}$. Therefore $A/\mathfrak{b}$ is $\mathfrak{F}$-discrete so that $\mathfrak{b}\in\mathfrak{F}$. Now, assume that $D$ is a closed subset of $\operatorname{Sp}A$. Then $\operatorname{Sp}A\setminus D$ is an open subset of $\operatorname{Sp}\mathcal{A}$; and hence $\mathcal{X}(D)=\operatorname{ZSupp}^{-1}(\operatorname{Sp}A\setminus D)$ is a localizing subcategory of $\operatorname{Mod}A$. It follows from section 3 and section 6 that $\operatorname{ZSupp}\mathcal{X}(D)=\operatorname{Sp}A\setminus D=\bigcup_{\lambda\in\varLambda}\operatorname{ZSupp}A/\mathfrak{a}_{\lambda}$ for some set $\varLambda$. The definition implies that $\mathfrak{F}_{D}=\{\mathfrak{a}_{\lambda}|\hskip 2.84544pt\lambda\in\varLambda\}$. To prove $D(\mathfrak{F}_{D})=D$, it is clear that $D\subseteq D(\mathfrak{F}_{D})$. Conversely, for any $E\in D(\mathfrak{F}_{D})$, we have $E\notin\operatorname{ZSupp}A/\mathfrak{a}_{\lambda}$ for every $\lambda\in\varLambda$. This implies that $E\notin\operatorname{ZSupp}\mathcal{X}_{D}$; and consequently $E\in D$. ∎
[…]
Let $A$ be a coherent ring. Then $\operatorname{ZASpec}\mathcal{A}$ is a retract subspace of $\operatorname{Zg}\mathcal{A}$.
Proof.
We define $\theta:\operatorname{Zg}A\rightarrow\operatorname{ZASpec}A$ given by $E\rightarrow E(A/{\mathfrak{p}_{E}})$. If $E(\alpha)\in\operatorname{ZASpec}\mathcal{A}$, then there exists a prime ideal $\mathfrak{p}=\mathfrak{p}_{E(\alpha)}$ of $A$ such that $\alpha=\overline{A/\mathfrak{p}}$ by section 5. Hence $E(\alpha)=E(A/\mathfrak{p})$ so that $\theta(E(\alpha))=E(\alpha)$. For every finitely presented $A$-module $M$, we show that $\theta^{-1}(\operatorname{ASupp}M)=\mathcal{O}(M)$. If $E\in\theta^{-1}(\operatorname{ASupp}M)$, we have $\theta(E)=E(A/\mathfrak{p}_{E})\in\operatorname{ASupp}M$ so that $E(A/\mathfrak{p}_{E})\in\mathcal{O}(M)$ by section 3. It now follows from [GP, Theorem 1.4] that $E\in\mathcal{O}(M)$. Conversely if $E\in\mathcal{O}(M)$, then $\theta(E)=E(A/\mathfrak{p}_{E})\in\operatorname{ASupp}M$ by [GP, Theorem 1.4]. Therefore $E\in\theta^{-1}(\operatorname{ASupp}M)$. ∎
[…]
Let $A$ be a coherent ring. Then =$\mathcal{F}\mathcal{I}=\{\operatorname{ASupp}A/\mathfrak{a}|\hskip 2.84544pt\mathfrak{a}$ is a finitely generated ideal of $A\}$ forms a basis for the topology $\operatorname{ZASpec}A$.
Proof.
For any finitely generated ideals $\mathfrak{a}$ and $\mathfrak{b}$ of $A$, similar to section 6, we have $\operatorname{ASupp}A/\mathfrak{a}\cap\operatorname{ASupp}A/\mathfrak{b}=\operatorname{ASupp}A/(\mathfrak{a}+\mathfrak{b})$. On the other hand, assume that $M$ is a finitely presented $A$-module. Then $M$ is finitely generated, say $M=<x_{1},\dots,x_{n}>$. Set $M_{0}=0$ and $M_{k}=<x_{1},\dots x_{k}>$ for $1\leq k\leq n$. Then $M_{i}/M_{i-1}$ is cyclic and finitely presented for each $i$. Thus there exist finitely generated ideals $\mathfrak{a}_{i}$ such that $M_{i}/M_{i-1}\cong A/\mathfrak{a}_{i}$ for all $1\leq i\leq n$. It is clear that $\operatorname{ASupp}M=\bigcup_{i=1}^{n}\operatorname{ASupp}A/\mathfrak{a}_{i}$. ∎
[…]
Let $A$ be a commutative coherent ring. Then the maps
@@
For every finitely generated ideal $\mathfrak{a}$ of $A$, it is known that $\operatorname{Spec}A\setminus V(\mathfrak{a})$ is a qausi-compcat open subset of $\operatorname{Spec}A$. Therefore for every open subset $\mathcal{V}$ of $\operatorname{Spec}^{\star}A$, we have $\mathcal{V}=\bigcup_{\gamma\in\varGamma}V(\mathfrak{a}_{\gamma})$ where $\mathfrak{a}_{\gamma}$ ranges over finitely generated ideals of $A$. By the same argument used in the proof of section 6, we have $\mathcal{O}_{V(\mathfrak{a}_{\gamma})}=\operatorname{ASupp}A/\mathfrak{a}_{\gamma}$ for all $\gamma\in\varGamma$; and hence $\{\overline{A/\mathfrak{p}}\in\operatorname{ZASpec}A|\hskip 2.84544pt\mathfrak{p}\in\mathcal{V}\}=\bigcup_{\gamma\in\varGamma}\operatorname{ASupp}A/\mathfrak{a}_{\gamma}$ is an open subset of $\operatorname{ZASpec}A$. Let $\mathcal{O}$ be an open subset of $\operatorname{ZASpec}\mathcal{A}$. By section 6, we have $\mathcal{O}=\bigcup_{\varLambda}\operatorname{ASupp}A/\mathfrak{a}_{\lambda}$, where $\mathfrak{a}_{\lambda}$ ranges over finitely generated ideals. Hence $\{\mathfrak{p}\in\operatorname{Spec}A|\hskip 2.84544pt\overline{A/\mathfrak{p}}\in\mathcal{O}\}=\bigcup_{\lambda\in\varLambda}V(\mathfrak{a}_{\lambda})$ is a open subset of $\operatorname{Spec}^{\star}A$. Now, it is easy to see that $\mathcal{O}_{\mathcal{V}_{\mathcal{O}}}=\mathcal{O}$ and $\mathcal{V}_{\mathcal{O}_{\mathcal{V}}}=\mathcal{V}$. ∎

### v2

@@ 5. In the case of commutative coherent rings
For every finitely generated ideal $\mathfrak{a}$ of $A$, it is known that $\operatorname{Spec}A\setminus V(\mathfrak{a})$ is a qausi-compact open subset of $\operatorname{Spec}A$. Therefore for every open subset $\mathcal{V}$ of $\operatorname{Spec}^{\star}A$, we have $\mathcal{V}=\bigcup_{\gamma\in\varGamma}V(\mathfrak{a}_{\gamma})$ where $\mathfrak{a}_{\gamma}$ ranges over finitely generated ideals of $A$. By the same argument used in the proof of section˜5, we have $\mathcal{O}_{V(\mathfrak{a}_{\gamma})}=\operatorname{ASupp}A/\mathfrak{a}_{\gamma}$ for all $\gamma\in\varGamma$; and hence $\{\overline{A/\mathfrak{p}}\in\operatorname{ZASpec}A|\hskip 2.84544pt\mathfrak{p}\in\mathcal{V}\}=\bigcup_{\gamma\in\varGamma}\operatorname{ASupp}A/\mathfrak{a}_{\gamma}$ is an open subset of $\operatorname{ZASpec}A$. Let $\mathcal{O}$ be an open subset of $\operatorname{ZASpec}\mathcal{A}$. By section˜5, we have $\mathcal{O}=\bigcup_{\varLambda}\operatorname{ASupp}A/\mathfrak{a}_{\lambda}$, where $\mathfrak{a}_{\lambda}$ ranges over finitely generated ideals. Hence $\{\mathfrak{p}\in\operatorname{Spec}A|\hskip 2.84544pt\overline{A/\mathfrak{p}}\in\mathcal{O}\}=\bigcup_{\lambda\in\varLambda}V(\mathfrak{a}_{\lambda})$ is a open subset of $\operatorname{Spec}^{\star}A$. Now, it is easy to see that $\mathcal{O}_{\mathcal{V}_{\mathcal{O}}}=\mathcal{O}$ and $\mathcal{V}_{\mathcal{O}_{\mathcal{V}}}=\mathcal{V}$. ∎
Lemma 5.4.
The set $\{\operatorname{\mathcal{O}}(A/\mathfrak{a})|\hskip 2.84544pt\mathfrak{a}$ is a finitely generated ideal of $\mathcal{A}\}$ forms a basis for the topology $\operatorname{Zg}A$.
Proof.
We show that $\operatorname{Zg}A=\bigcup_{\mathfrak{a}\in\text{fg-}A}\mathcal{O}(A/\mathfrak{a})$. Any indecomposable injective module $E$ contains a cyclic submodule $A/\mathfrak{c}$. This implies that $E\in\operatorname{\mathcal{O}}(A/\mathfrak{c})$. If $\mathfrak{c}=0$, there in nothing to prove. Otherwise, $\mathfrak{c}$ contains a finitely generated ideal $\mathfrak{d}$ so that $E\in\mathcal{O}(A/\mathfrak{d})$. The rest of the proof is similar to that in section˜5. ∎
A set $\mathfrak{F}$ of ideals of $A$ is said to be a Gabriel topology on $A$ if it satisfies:
(2). If $\mathfrak{a}$, $\mathfrak{b}$ belong to $\mathfrak{F}$, then $\mathfrak{a}\cap\mathfrak{b}\in\mathfrak{F}$.
A Gabriel topology $\mathfrak{F}$ is said to have a basis of finitely generated ideals, if any ideal $\mathfrak{a}\in\mathfrak{F}$ contains a finitly generated ideal $\mathfrak{b}\in\mathfrak{F}$. Let $E$ be an indecomposable injective module in $\operatorname{Zg}A$. Clearly, the Gabriel topology
$\mathfrak{F}_{E}=\{\mathfrak{a}|\hskip 2.84544ptE\notin\operatorname{\mathcal{O}}(A/\mathfrak{a})\}$
is associated to $\{M\in\operatorname{Mod}A|\hskip 2.84544pt\operatorname{Hom}_{A}(M,E)=0\}$, the localizing subcategory cogenerated by $E$. More generally, let $\mathcal{E}$ be a subset of $\operatorname{Zg}A$. To $\mathcal{E}$, we associated a Gabriel topology
$\mathfrak{F}_{\mathcal{E}}=\bigcap_{E\in\mathcal{E}}\mathfrak{F}_{E}=\{\mathfrak{a}|\hskip 2.84544pt\operatorname{\mathcal{O}}(A/\mathfrak{a})\cap\mathcal{E}=\emptyset\}.$
$D(\mathfrak{F})=\{E\in\operatorname{Zg}A|\hskip 2.84544ptE\notin\operatorname{\mathcal{O}}(A/\mathfrak{a})\hskip 2.84544pt{\rm for\hskip 2.84544ptevery}\hskip 2.84544pt\mathfrak{a}\in\mathfrak{F}\}=\operatorname{Zg}A\setminus\bigcup_{\mathfrak{a}\in\mathfrak{F}}\operatorname{\mathcal{O}}(A/\mathfrak{a}).$
Given a Gabriel topology $\mathfrak{F}$, an $A$-module $M$ is said to be $\mathfrak{F}$-discrete if $\operatorname{Ann}(x)\in\mathfrak{F}$ for all $x\in M$. The following theorem provides a new classification of Gabriel topologies on $A$.
Lemma 5.5.
Let $D$ be a closed subset of $\operatorname{Zg}A$. Then $\mathfrak{F}_{D}$ has a basis of finitely generated ideals of $A$.
Proof.
We observe that $\operatorname{Zg}A\setminus D$ is an open subset of $\operatorname{Zg}A$; and hence by section˜5, there exists a set of finitely generated ideals $\{\mathfrak{b}_{\lambda}|\hskip 2.84544pt\lambda\in\varLambda\}\subseteq\mathfrak{F}_{D}$ such that $\operatorname{Zg}A\setminus D=\bigcup_{\lambda\in\varLambda}\mathcal{O}(A/\mathfrak{b}_{\lambda})$. Now, given an ideal $\mathfrak{a}\in\mathfrak{F}_{D}$, by the definition, $\mathcal{O}(A/\mathfrak{a})\subseteq\bigcup_{\lambda\in\varLambda}\mathcal{O}(A/\mathfrak{b}_{\lambda})$ so that $A/\mathfrak{a}$ is a quotient of $A/\mathfrak{c}\in\sqrt{\langle A/\mathfrak{b}_{\lambda}|\hskip 2.84544pt\lambda\in\varLambda\rangle}$ by section˜4. This implies that $\mathfrak{c}\subseteq\mathfrak{a}$ and there exists a finite filtration of submodules of $A/\mathfrak{c}$
$0=X_{0}\subset X_{1}\subset\dots\subset X_{n-1}\subset X_{n}=A/\mathfrak{c}$
such that each $X_{i}/X_{i-1}$ is a subquotient of $A/\mathfrak{b}_{\lambda_{i}}$ for some $\lambda_{i}\in\varLambda$. Setting $\mathfrak{b}=\mathfrak{b}_{\lambda_{1}}\mathfrak{b}_{\lambda_{2}}\dots\mathfrak{b}_{\lambda_{n}}$, it follows that $\mathfrak{b}\subset\mathfrak{c}$ and $\mathfrak{b}\in\mathfrak{F}_{D}$ by [St, Chap VI, p. 147, Lemma 5.3]. ∎
There are many classifications of Gabriel topologies on $A$ (e.g., see [G, St]). For commutative coherent rings, the following theorem provides a new classification of Gabriel topologies on $A$. The set of finitely generated modules is denoted by fg-$A$.
Theorem 5.6.
There are bijective correspondence between:
$\rm(1)$} the lattice of open subsets of $\operatorname{ZASpec}A$,
$\rm(2)$} the lattice of closed subsets of $\operatorname{Zg}A$,
$\rm(3)$} the lattice of Gabriel topologies having bases of finitely generated ideals of $A$,
$\rm(4)$} The lattice of localizing subcategory of finite type of $\operatorname{Mod}A$.
These bijections are defined as follows:
${\rm(1)}\longleftrightarrow{\rm(2)}$: $\operatorname{ASupp}A/\mathfrak{a}\longleftrightarrow\operatorname{Zg}A\setminus\mathcal{O}(A/\mathfrak{a})$ for any finitely generated ideal $\mathfrak{a}$.
${\rm(2)}\longleftrightarrow{\rm(3)}$: $D\stackrel{{\scriptstyle\mathfrak{F}}}{{\mapsto}}\mathfrak{F}_{D}$ and $\mathfrak{F}\stackrel{{\scriptstyle D}}{{\mapsto}}D(\mathfrak{F})$,
${\rm(3)}\longleftrightarrow{\rm(4)}$: $\mathfrak{F}\mapsto\mathcal{X}_{\mathfrak{F}}=\overrightarrow{\sqrt{\langle A/\mathfrak{a}|\hskip 2.84544pt\mathfrak{a}\in\mathfrak{F}\cap\operatorname{fg-}A\rangle}}$ and $\mathcal{X}\mapsto\mathfrak{F}_{\mathcal{X}}=\{\mathfrak{a}|\hskip 2.84544ptA/\mathfrak{a}\in\mathcal{X}\}$.
Proof.
${\rm(1)}\longleftrightarrow{\rm(2)}$ follows from section˜3, section˜4, section˜5 and section˜5.
${\rm(2)}\longleftrightarrow{\rm(3)}$. Let $\mathfrak{F}$ be a Gabriel topology on $A$ having a basis of finitely generated ideal of $A$ and $D$ be a closed subset of $\operatorname{Zg}A$. Then section˜5 implies that $\mathfrak{F}_{D}$ has a basis of finitely generated ideals and $D(\mathfrak{F})=\operatorname{Zg}A\setminus\bigcup_{\mathfrak{a}\in\mathfrak{F}}\operatorname{\mathcal{O}}(A/\mathfrak{a})=\operatorname{Zg}A\setminus\bigcup_{\mathfrak{a}\in\mathfrak{F}\cap\operatorname{fg-}A}\operatorname{\mathcal{O}}(A/\mathfrak{a})$ is a closed subset of $\operatorname{Zg}A$ by section˜5. We show that $\mathfrak{F}_{D(\mathfrak{F})}=\mathfrak{F}$. Given $\mathfrak{a}\in\mathfrak{F}$, the definition implies that $\operatorname{Hom}_{A}(A/\mathfrak{a},D(\mathfrak{F}))=0$ so that $\mathfrak{a}\in\mathfrak{F}_{D(\mathfrak{F})}$. Conversely, if $\mathfrak{b}\in\mathfrak{F}_{D(\mathfrak{F})}$, then $\operatorname{Hom}_{A}(A/\mathfrak{b},D(\mathfrak{F}))=0$ so that $\operatorname{\mathcal{O}}(A/\mathfrak{b})\subseteq\bigcup_{\mathfrak{a}\in\mathfrak{F}}\operatorname{\mathcal{O}}(A/\mathfrak{a})$. It then follows from section˜5 that $\operatorname{\mathcal{O}}(A/\mathfrak{b})\subseteq\bigcup_{\mathfrak{a}\in\mathfrak{F}\cap{\text{fg-}}A}\operatorname{\mathcal{O}}(A/\mathfrak{a})=\mathcal{O}(\overrightarrow{\sqrt{\langle A/\mathfrak{a}|\hskip 2.84544pt\mathfrak{a}\in\mathfrak{F}\cap{\text{fg-}}A\rangle}})$. Now, section˜4 implies that $A/\mathfrak{b}\in\overrightarrow{\sqrt{\langle A/\mathfrak{a}|\hskip 2.84544pt\mathfrak{a}\in\mathfrak{F}\cap{\text{fg-}}A\rangle}}$. Therefore, $A/\mathfrak{b}$ is a quotient of some $A/\mathfrak{a}$ for $\mathfrak{a}\in\mathfrak{F}\cap{\text{fg-}}A$. This implies that $\mathfrak{a}\subseteq\mathfrak{b}$ so that $\mathfrak{b}\in\mathfrak{F}$. To prove $D(\mathfrak{F}_{D})=D$, since $\operatorname{Zg}A\setminus D$ is an open subset of $\operatorname{Zg}\mathcal{A}$, by section˜5, there exists a set of finitely generated ideals $\{\mathfrak{a}_{\lambda}|\hskip 2.84544pt\lambda\in\varLambda\}$ of $A$ such that $\operatorname{Zg}A\setminus D=\bigcup_{\lambda\in\varLambda}\operatorname{\mathcal{O}}(A/\mathfrak{a}_{\lambda})$. The inclusion $D\subseteq D(\mathfrak{F}_{D})$ is clear. Conversely, for any $E\in D(\mathfrak{F}_{D})$, since $\mathfrak{a}_{\lambda}\in\mathfrak{F}_{D}$ for any $\lambda\in\varLambda$, we have $E\notin\operatorname{\mathcal{O}}(A/\mathfrak{a}_{\lambda})$ for any $\lambda\in\varLambda$. Consequently $E\in D$.
${\rm(3)}\longleftrightarrow{\rm(4)}$. Given a Gabriel topology $\mathfrak{F}$ with a basis of finitely generated ideals of $A$, it follows from [Kr, Theorem 2.8] that $\mathcal{X}_{\mathfrak{F}}$ is a localizing subcategory of finite type of $\mathcal{A}$. Moreover, for a localizing subcategory $\mathcal{X}$ of $\operatorname{Mod}A$ of finite type, it follows from [St, Chap VI, p. 146, Theorem 5.1] that $\mathfrak{F}_{\mathcal{X}}$ is a Gabriel topology on $A$. If $\mathfrak{a}\in\mathfrak{F}_{\mathcal{X}}$, then we have $A/\mathfrak{a}=\underset{\rightarrow}{\rm lim}M_{i}$ is the direct limit of objects $M_{i}$ in $\mathcal{X}\cap\operatorname{fp-}A$. Then $\mathfrak{a}$ contains a finitely generated ideal $\mathfrak{b}$ such that $A/\mathfrak{b}\in\mathcal{X}$. Hence $\mathfrak{F}_{\mathcal{X}}$ has a basis of finitely generated ideals of $A$. We first show that $\mathfrak{F}_{\mathcal{X}_{\mathfrak{F}}}=\mathfrak{F}$. Given $\mathfrak{a}\in\mathfrak{F}$, it contains a finitely generated ideal $\mathfrak{b}\in\mathfrak{F}$ so that $A/\mathfrak{b}\in\mathcal{X}$. Then $A/\mathfrak{a}\in\mathcal{X}_{\mathfrak{F}}$; and hence $\mathfrak{a}\in\mathfrak{F}_{\mathcal{X}_{\mathfrak{F}}}$. Conversely, if $\mathfrak{a}\in\mathfrak{F}_{\mathcal{X}_{\mathfrak{F}}}$, then $A/\mathfrak{a}\in\mathcal{X}_{\mathfrak{F}}$. By a similar argument as mentioned above, $\mathfrak{a}$ contains a finitely generated ideal $\mathfrak{b}\in\mathfrak{F}$ so that $\mathfrak{a}\in\mathfrak{F}$. Now, we show that $\mathcal{X}_{\mathfrak{F}_{\mathcal{X}}}=\mathcal{X}$. Given $M\in\mathcal{X}$, since $\mathcal{X}$ and $\mathcal{X}_{\mathfrak{F}_{\mathcal{X}}}$ are localizing subcategory of finite type of $\mathcal{A}$, we may assume that $M$ is finitely presented. Assume that $M=<x_{1},\dots,x_{n}>$, $M_{0}=0$ and $M_{k}=<x_{1},\dots x_{k}>$ for $1\leq k\leq n$. Then $M_{i}/M_{i-1}=A/\mathfrak{a}_{i}\in\mathcal{X}$ for $1\leq i\leq n$. Hence $M_{i}/M_{i-1}\in\mathcal{X}_{\mathfrak{F}_{\mathcal{X}}}$ for $1\leq i\leq n$; and consequently $M\in\mathcal{X}_{\mathfrak{F}_{\mathcal{X}}}$. ∎